# Using iterated local alignment to aggregate trajectory data into a traffic flow map

### v5

@@ 2.1 Map matching
1:Input: $G$ trajectory, $M$ map matching API, $R$ route finding API, ${\bm{n}}_{W}$waypoints
2:Output: $R(G)$ map matched route
3:Compute initial map matched route $M(G)$ from empirical trajectory $G$
4:for $i:=1$ to $w$ do
5:     Set $n_{W}:=n_{W,i}$
6:     Sample waypoints ${\bm{w}}_{1},\dots,{\bm{w}}_{n_{W}}$ from initial map matched $M(G)$
7:     Compute map matched routes $R_{i}(G):=R({\bm{w}}_{1},\dots,{\bm{w}}_{n_{W}})$
8:Select minimal route $R(G):=\operatorname{argmin}_{i\in\{1,\dots,w\}}\,\operatorname{DTW}(R_{i}(G),G)$
We set $R$ to be the Valhalla Odin turn-by-turn route finding API to be consistent with our choice of $M$ as the Valhalla Meili map matching API. We set the number of waypoints as ${\bm{n}}_{W}=3,13,23,33,43,63,83$ since these ${\bm{n}}_{W}$ cover a range of possible values which likely includes an approximately optimal value. The map matched Hannover trajectories are displayed as the blue lines in Figure 1.
2.2 Flow aggregation
To the road network graph $\mathcal{N}=\mathcal{N}(\mathcal{E},\mathcal{V})$, we add the traffic flows on each of the network edges to form a flow map =$\mathcal{F}=\{(f,{\bm{f}}):{\bm{f}}\in\mathcal{N},f>0\}=\{(f_{1},{\bm{f}}_{1}),\dots,(f_{n_{\mathcal{F}}},{\bm{f}}_{n_{\mathcal{F}}})\}$ composed of $n_{\mathcal{F}}$ flows, where each ${\bm{f}}_{i}$ is composed of edges in $\mathcal{E}$, with traffic flow $f_{i}$, $i=1,\dots,n_{\mathcal{F}}$. Our goal is to estimate the traffic flows $f_{1},\dots,f_{n_{\mathcal{F}}}$ and the flow network graph $\mathcal{F}$. For exactly aligned trajectories, to compute the flow aggregation, we deploy the overline function in the R package stplanr, which we refer to as ST_OVERLINE_PLANR (Lovelace and Ellison, 2018). That is, starting with the map matched routes =$\mathcal{R}=\{R(G_{1}),\dots,R(G_{n})\}$, the flow map is $\mathcal{F}={\tt ST\_OVERLINE\_PLANR}(\mathcal{R})$.
To compute an accurate flow map from noisy, empirical trajectories relies on resolving the crucial problem of how to aggregate similar, but not exactly overlapping, road segments. Many solutions have been offered, such as edge bundling (Zhou et al., 2013) and rasterisation (Wood et al., 2010, Morgan and Lovelace, 2021). Edge bundling consists of clustering trajectory linestrings and replacing all cluster members with a single representative linestring. These (and subsequent) authors conclude that it performs poorly when applied to noisy trajectories at the road level, and remains mostly suited to large-scale aggregations, such as desire line maps (Tobler, 1987). We do not consider edge bundling further since it is not suitable for our purposes. Rasterisation relies on converting the flow map into a raster matrix, and aggregating the flows within the same raster pixel neighbourhood. Whilst rasterisation is able to improve flow aggregation, it depends highly on the raster pixel neighbourhood size, and leads to a loss of spatial resolution.
We propose an alternative aggregation which solves both of these problems. We achieve this goal by local internal alignment between the map matched routes. By internal alignment, we mean that we align the routes with each other, rather than to the external road network graph. By local alignment, we mean that we align segments of the routes, rather than the complete routes. Our proposal is divided into several algorithms, which we develop within the R statistical analysis environment due to its integrated access to advanced statistical and geospatial analysis methods.
2.3 Node snapping and node splitting
The proposed algorithm is ST_SNAPNODE, where the boundary points of the traffic flow linestrings are snapped to each other. Since the former are also nodes of the flow map, this gives the name to the algorithm. We focus on snapping these nodes since the linestring misalignments are in part caused by the existence of nodes which are close to each other but not exactly equal.
Since we are searching for points which are close to together, then this is well-suited to statistical clustering. There are many statistical clustering algorithms available, and we focus on hierarchical clustering. To resolve the computational bottleneck of hierarchical clustering, we approximate the 1-pass complete linkage clustering by a nested 2-pass clustering. We begin with an efficient single linkage clustering in the R package fastcluster (Müllner, 2013). Since single linkage can result in chain-like clusters, we compute a subsequent complete linkage clustering to break these potential chains. In this nested approach, the complete linkage distance matrix is calculated only within each single linkage cluster, and so it is less likely to reach computational limits.
@@ 2.3 Node snapping and node splitting
Algorithm 2 is a description of ST_SNAPNODE. The inputs are the $n_{\mathcal{F}}$ flow linestrings =$\mathcal{F}=\{(f_{1},{\bm{f}}_{1}),\dots,(f_{n_{\mathcal{F}}},{\bm{f}}_{n_{\mathcal{F}}})\}$ and the snap tolerance $\varepsilon_{S}$. As its name suggests, all node points within $\varepsilon_{S}$ distance of each other are snapped together. In Step 1, we extract the boundary points of all flow linestrings. In Steps 2–3, we compute a single linkage clustering on all boundary points, and cut the dendrogram at height $\varepsilon_{S}$, resulting in $C^{\prime}$ clusters. In Steps 4–7, within each of these single linkage clusters, we compute a complete linkage clustering, and cut the dendrogram at height $\varepsilon_{S}$. This divides each single linkage cluster into $C^{\prime\prime}$ clusters, where all cluster members are at most $\varepsilon_{S}$ distance from each other. In Steps 8–11, we compute the weighted spatial centroid in each of the $C^{\prime\prime}$ complete linkage clusters, weighted by the traffic flow values. In Step 12, we renumber the cluster labels from the above nested clusterings to approximate a 1-pass complete linkage clustering into $C^{*}$ clusters. In Steps 13–16, within each of these $C^{*}$ clusters, for all points of the corresponding flow linestrings which are closer than $\varepsilon_{S}$ distance to the centroid, we snap them to the centroid. In Steps 17–18, we collate and sort all the $n_{\mathcal{F}^{*}}$ snapped linestrings =$\mathcal{F}^{*}=\{(f_{1}^{*},{\bm{f}}_{1}^{*}),\dots,(f_{n_{\mathcal{F}^{*}}}^{*},{\bm{f}}_{n_{\mathcal{F}^{*}}}^{*})\}$. In comparison to the unsnapped linestrings $\mathcal{F}$, there are fewer snapped linestrings $\mathcal{F}^{*}$ which have higher flow values, are more interconnected, and share more exactly overlapping sub-segments.
@@
19:Collate snapped linestrings $\{\mathcal{F}_{1}^{*},\dots\mathcal{F}_{C^{\prime*}}^{*}\}$ into :=$\mathcal{F}^{*}:=\{(f_{1}^{*},{\bm{f}}_{1}^{*}),\dots,(f_{n_{\mathcal{F}^{*}}}^{*},{\bm{f}}_{n_{\mathcal{F}^{*}}}^{*})\}$
@@
Node snapping is focused near the boundary points of the flow linestrings, so it does not address misalignments far from the boundary points. Moreover there remain some linestrings whose intersection is not correctly computed by ST_OVERLINE_PLANR. The solution to both of these problems is the explicit addition of the missing intersection nodes to these linestrings.
ST_SPLITNODE is the collation of a couple of algorithms which explicitly add the nodes at the intersections in the interior of linestrings, and splits these linestrings into simple linestring segments at these added nodes. This gives the name to this method. The inputs to Algorithm 3 are the flow linestrings $\mathcal{F}$ and the node splitting type $S$. Our method deploys for $S=$ “subdivision” in Steps 1–3, to_spatial_subdivision in the sfnetworks package (van der Meer et al., 2023), and for $S=$ “unary” in Steps 4–5, geos_unary_union in the geos package (Dunnington and Pebesma, 2023). The first option tends to find fewer intersection nodes and is similar to the intersection computation performed by ST_OVERLINE_PLANR. The second option finds more missing intersection nodes, and leads to more comprehensive flow aggregation. We require both types of node splitting to compute a locally aligned flow map.
@@
7:     $\mathcal{F}^{*}:=$ geos::geos_unary_union$(\mathcal{N}^{*})$
In Figure 2(a) are the flow segments without any node splitting or node snapping. There are missing intersection nodes, and there are intersection nodes which are close to each other. In Figure 2(b) are the flow segments after node splitting ($S=$ “unary”) and then node snapping (with snap tolerance $\varepsilon_{S}=4$ m). The result is that there are fewer flow segments with higher flow values. Due to the combined action of ST_SPLITNODE and ST_SNAPNODE, Figure 2(b) is an improvement over Figure 2(a).
Refer to caption Refer to caption
Figure 2: Unsplit/unsnapped and node split/node snapped flow segments, with snap tolerance $\varepsilon_{S}=4$ m. (a) Unsplit and unsnapped. (b) Node split and node snapped.
2.4 Line blending to align similar linestrings with local reference
Let ${\bm{f}}_{\rm ref}$ be a reference linestring from $\mathcal{F}$. The set of $n_{\mathcal{F}_{\rm cand}}$ linestrings from \$\mathcal{F}\backslash\{{\bm{f}}_{\rm ref}\}$ which satisfy Equation (1) are the candidate linestrings $\mathcal{F}_{\rm cand}=\{{\bm{f}}_{{\rm cand},1},\dots,{\bm{f}}_{{\rm cand},n_{\mathcal{F}_{\rm cand}}}\}$, with the convention that if $n_{\mathcal{F}_{\rm cand}}=0$ then $\mathcal{F}_{\rm cand}$ is the empty set. We call our approach “line blending”, since we blend the candidate linestrings onto the reference linestring. The inputs in Algorithm 4 are the reference linestring ${\bm{f}}_{\rm ref}$, the $n_{\mathcal{F}_{\rm cand}}$ candidate linestrings $\mathcal{F}_{\rm cand}$, and the blend tolerance $\varepsilon$. The output are the modified reference linestring and $n_{\mathcal{F}_{\rm cand}}$ modified candidate flow linestrings, all with added interior points for accurate calculation of exactly equal linestring segments. In Step 1, we initialize a local network graph $\mathcal{N}^{*}$ with the reference linestring. In Steps 2–3, we extract all points of the candidate linestrings, and use the network_blend function in the sfnetworks package (which we denote as ST_NETWORK_BLEND) to blend efficiently these points into the reference linestring (van der Meer et al., 2023). ST_NETWORK_BLEND projects the candidate linestrings onto the reference linestring, and explicitly adds them to the network, thereby creating new edges in $\mathcal{N}^{*}$. The result is a local network graph with more, shorter edges and with nodes at the projected candidate points, and whose union is the reference linestring. In Steps 4–5, we extract the $n_{\mathcal{F}_{\rm cand}}$ blended candidate linestrings with these added interior points by applying the shortest path search between the start and end point of each blended candidate linestring along the network graph $\mathcal{N}^{*}$. Step 6 involves collating the blended candidate linestrings. Step 7 is the equivalent for the blended reference linestring.
@@ 2.4 Line blending to align similar linestrings with local reference
9:${\bm{f}}^{*}_{{\rm ref}}$ := shortest path from $\operatorname{Start}({\bm{f}}_{{\rm ref}})$ to $\operatorname{End}({\bm{f}}_{{\rm ref}})$ along network $\mathcal{N}^{*}$
The result from ${\tt ST\_LINEBLEND}$ is the modified reference linestring and the projected candidate linestrings with added interior points. If there are many candidate linestrings, then this may lead to many sub-segments in the projected linestrings, each with their own flow values. So we assign the rounded value of the weighted mean flow, weighted by the sub-segment length, to all sub-segments.
In Figure 3(a), we illustrate this line blending with the flow linestrings $(7,ABC)$ in blue and $(2,AD)$ in orange. As these two linestrings do not share a common sub-segment, so flow aggregation does not modify them. Since $AD\subset{\tt ST\_BUFFER}(ABC,4\,\mathrm{m})$ (the pale blue rectangle), and $2<7$, then $ABC$ is the reference ${\bm{f}}_{\rm ref}$ and $AD$ is the candidate linestring ${\bm{f}}_{\rm cand}$, according to Equation (1). Line blending involves projecting $AD$ to $ABC$, so $D$ is projected to $D^{\prime}$ (which lies exactly on the reference linestring), and $B$ is added explicitly to this projected linestring. We also add $D^{\prime}$ to the reference linestring. The reference linestring becomes ${\bm{f}}^{*}_{\rm ref}=(7,ABD^{\prime}C)$, and the candidate ${\bm{f}}^{*}_{\rm cand}=(2,ABD^{\prime})$. Now ${\bm{f}}^{*}_{\rm ref},{\bm{f}}^{*}_{\rm cand}$ share the sub-segment $ABD^{\prime}$ exactly, and ST_OVERLINE_PLANR gives the aggregated flows as $(9,ABD^{\prime})$ and $(2,D^{\prime}C)$. We take the rounded value of the weighted mean of these two flow linestrings. The result is a single linestring $(9,ABD^{\prime}C)$, as shown in Figure 3(b).
Refer to caption Refer to caption
Figure 3: Line blending leads to correct flow aggregation. (a) Before line blending and flow aggregation. Reference linestring $(7,ABC)$ in blue, candidate linestring $(2,AD)$ in orange. (b) After blending candidate into reference linestring, with tolerance $\varepsilon=4$ m, and flow aggregation. Blended linestring is $(9,ABD^{\prime}C)$.
After aggregation, the projected point $D^{\prime}$ in $ABD^{\prime}C$ appears to be redundant. Suppose that we have another linestring $(3,ABC)$ without $D^{\prime}$. Then ST_OVERLINE_PLANR does not aggregate $(9,ABD^{\prime}C)$ with $(3,ABC)$ even if $D^{\prime}$ is exactly coincident on the segment $BC$, since this algorithm does not explicitly add any missing intersection nodes prior to aggregation (Morgan and Lovelace, 2021). It outputs three linestrings $(12,AB),(9,BC),(3,BD^{\prime}C)$, rather than a single linestring $(12,ABD^{\prime}C)$. So we retain these intermediate points in the flow linestrings to assist in subsequent flow aggregation.
More complex situations arise when other linestrings touch the candidate linestring, but are not candidates themselves for blending. Since line blending projects the candidates to the reference linestring, then we have to also project these other linestrings to avoid leaving a gap in the updated flow map. This procedure ST_SNAP_CAND_TOUCH is outlined in Algorithm 5. The inputs are the reference linestring ${\bm{f}}_{\rm ref}$, the $n_{\mathcal{F}_{\rm cand}}$ candidate linestrings $\mathcal{F}_{\rm cand}$, the $n_{\mathcal{F}_{\rm ct}}$ candidate-touching linestrings $\mathcal{F}_{\rm ct}$, and the snap tolerance $\varepsilon_{S}$. In Steps 1–8, we iterate over each candidate-touching linestring. In Steps 2–3, we compute the intersections between the candidate-touching linestring and the boundary points of the candidate linestrings, and the respective distances. In Steps 4–7, if this intersection point is within $\varepsilon_{S}$ distance to the boundary points of ${\bm{f}}_{\rm ref}$, then we snap the candidate-touching linestring to the closest boundary point. In Step 8, if the intersection point is not within $\varepsilon_{S}$ distance, then we snap it to the nearest point on ${\bm{f}}_{\rm ref}$. These snapping operations are similar to those in Steps 15–16 in ST_SNAPNODE in Algorithm 2, and ensure that we maintain connectivity between $\mathcal{F}_{\rm ct}$ and ${\bm{f}}_{\rm ref}$. In Step 9, we collate these snapped linestrings.
Algorithm 5 ST_SNAP_CAND_TOUCH – Snap candidate-touching linestrings onto reference
@@
Figure 4(a) is an illustration of ST_SNAP_CAND_TOUCH with the reference $(7,ABC)$ (blue), the candidate linestring $(2,AD)$ (orange), and the blending buffer zone with blend tolerance $\varepsilon=4$ m (light blue). The candidate-touching linestring ${\bm{f}}_{\rm ct}=(5,DE)$ (purple) touches the candidate $AD$ at $D$. The orange line is entirely within the pale blue buffer zone, whereas the purple line extends outside of the buffer zone and so is not a candidate for blending to the reference linestring. If we apply ST_LINEBLEND to blend the candidate linestring, then $D$ is projected to $D^{\prime}$ to lie on the reference linestring, i.e., ${\bm{f}}_{\rm cand}^{*}=(2,ABD^{\prime})$. If we apply ST_SNAP_CAND_TOUCH with snap tolerance $\varepsilon_{S}=4$ m to the candidate-touching linestring, then the intersection point $D$ is projected to the boundary point $C$, i.e., ${\bm{f}}^{*}_{\rm ct}=(5,CE)$. Thus line blending, candidate-touching snapping, and flow aggregation yields that ${\bm{f}}^{*}_{\rm ct}=(5,CE)$ remains connected to ${\bm{f}}_{\rm ref}^{*}=(9,ABD^{\prime}C)$.
Refer to caption Refer to caption
Figure 4: Snap candidate-touching linestrings to reference linestring after line blending, with blend tolerance $\varepsilon=4$ m, snap tolerance $\varepsilon_{S}=4$ m. (a) Before line blending and snapping. Reference linestring $(7,ABC)$ in blue, candidate linestring $(2,AD)$ orange, and candidate touching linestring $(5,DE)$ purple. (b) After line blending and snapping. Reference linestring becomes $(9,ABD^{\prime}C)$, and candidate-touching linestring $(5,CE)$.
We generalise the reference linestring to be a sequence of $k$ connected edges, each with their own flow. If we treat this edge sequence momentarily as a single linestring, with flow equal to the weighted mean flow, weighted by the length of the $k$ edges, then we can apply Equation (1) to search for potential candidate linestrings. Due to computational limitations, we retain that candidates be single edges with single flow values. The generalisation to $k$-edge reference linestrings allows us to blend candidate linestrings which exceed the buffer zones of 1-edge reference linestrings.
We have described blending candidate and candidate-touching linestrings to a single reference linestring (possibly composed of $k$ edges). The next step is to determine the priority for line blending for a set of reference linestrings. We require that a reference linestring cannot be a candidate linestring to another reference linestring, and a candidate linestring is a candidate for one reference linestring only. These ensure that if two linestrings ${\bm{f}}_{2}\subseteq{\tt ST\_BUFFER}({\bm{f}}_{1},\varepsilon)$ and ${\bm{f}}_{1}\subseteq{\tt ST\_BUFFER}({\bm{f}}_{2},\varepsilon)$, then we select only one of them. The flow condition $f_{2}\leq f_{1}$ usually distinguishes the reference linestring, except when $f_{1}=f_{2}$. In this case, then we designate the longer linestring to be the reference.
@@
The inputs to ST_LINEBLEND_PRIORITY in Algorithm 6 are the flow linestrings $\mathcal{F}$, the number of connected edges for the reference linestrings $k$, and the blend tolerance $\varepsilon$. The output is a set of non-overlapping reference linestrings $\mathcal{F}_{\rm ref}$, a set of unique candidate linestrings $\mathcal{F}_{\rm cand}$, and a set of (potentially repeated) candidate-touching $\mathcal{F}_{\rm ct}$ linestrings. In Steps 1–2, we set up a network graph from the linestrings $\mathcal{F}$, and then extract $\mathcal{K}$, all simple paths (i.e., all paths composed of connected, unique edges) with $k$-edges. In Steps 3–4, we compute the weighted mean of the $k$ flow values, weighted by the length of individual edges, and assign this to the entire $k$-edge path. The function ${\rm Edge}(\cdot)$ extracts the component edges from a $k$-edge path. In Step 5, we sort the flow linestrings, in descending order of their flow values and length. In Steps 6–8, we construct a maximal set of non-overlapping $k$-edge paths $\mathcal{K}^{*}$. We initialise $\mathcal{K}^{*}$ to contain the first path. We iterate through $\mathcal{K}$ and, if this path in $\mathcal{K}$ does not overlap the current $\mathcal{K}^{*}$, then we add it to $\mathcal{K}^{*}$. For $k=1$, we bypass Steps 2–8. In Steps 9–18, we step through the ${\bm{k}}_{i}^{*}\in\mathcal{K}^{*}$, starting from the linestring with the highest flow and length. In Steps 11–13, if the current flow linestring ${\bm{k}}_{i}^{*}$ has no incident edges in the candidate linestrings $\mathcal{F}_{\rm cand}$, then we set it to be a reference linestring. In Step 13, we search for candidate linestrings for ${\bm{k}}_{i}^{*}$, from those linestrings which are not already reference or candidate linestrings, according to Equation (1). In Steps 14–17, if there is at least one candidate linestring, then we update $\mathcal{F}_{{\rm ref}},\mathcal{F}_{\rm cand}$, and $\mathcal{F}_{\rm ct}$. In Step 18, we extract the candidate $\mathcal{F}_{\rm cand}$ and candidate-touching $\mathcal{F}_{\rm ct}$ linestring sets.
@@
1:Input: $\mathcal{F}$ flow linestrings, $k$edges, $\varepsilon$ blend tolerance
@@
The local alignment of flow segments, in Algorithm 7, is established by combining Algorithms 26. The inputs to ST_OVERLINE_LINEBLEND are the flow linestrings $\mathcal{F}$, the split node type $S$, the number of connected edges $k$, the blend tolerance $\varepsilon$, and the snap tolerance $\varepsilon_{S}$. In Step 1, we node split the flow linestrings $\mathcal{F}$. In Step 2, we node snap the flow linestrings. In Step 3, we set up the priority for the line blending, and in Step 4 we store all the linestrings that will not be modified by the line blending in $\mathcal{F}^{c}$. In Steps 5–8, we iterate over each reference linestring. In Step 6, we update the reference and candidate flow linestrings by blending the candidate linestrings. In Step 7, we apply ST_OVERLINE_PLANR within each of the $k$-edges of the updated reference linestrings. In Step 8, we compute the weighted mean flow, weighted by the length of the corresponding sub-segments, and assign it as the flow value to all sub-segments. The result from Steps 6–8 is a modified $k$-edge reference linestring with $k$ aggregated flows. In Step 9, we snap the candidate-touching linestrings to the reference linestring to maintain connectivity. In Steps 10–11, we collate the modified linestrings from Steps 5–9, with the unmodified linestrings $\mathcal{F}^{c}$ from Step 4, and sort them in descending flow and length.

### v6

@@ A.1 Map matched route
1:Input: $G$ trajectory, $M$ map matching API
2:Output: $M(G)$ map matched route
3:Compute map matched route $M(G)$ from empirical trajectory $G$
A.2 Snap node clustering of linestrings
@@
The inputs to ST_SNAPNODE are the $n_{\mathcal{F}}$ flow linestrings $\mathcal{F}=\{(f_{1},{\bm{f}}_{1}),\dots,(f_{n_{\mathcal{F}}},{\bm{f}}_{n_{\mathcal{F}}})\}$ and the snap tolerance $\varepsilon_{S}$. The output are the snapped linestrings $\mathcal{F}^{*}$. In Step 1, we extract the boundary points of all flow linestrings. In Steps 2–3, we compute a single linkage clustering on all boundary points, and cut the dendrogram at height $\varepsilon_{S}$, resulting in $C^{\prime}$ clusters. In Steps 4–7, within each of these single linkage clusters, we compute a complete linkage clustering, and cut the dendrogram at height $\varepsilon_{S}$. This divides each single linkage cluster into $C^{\prime\prime}$ clusters, where all cluster members are at most $\varepsilon_{S}$ distance from each other. In Steps 8–11, we compute the weighted spatial centroid in each of the $C^{\prime\prime}$ complete linkage clusters, weighted by the traffic flow values. In Step 12, we renumber the cluster labels from the above nested clusterings to approximate a 1-pass complete linkage clustering into $C^{*}$ clusters. In Steps 13–16, within each of these $C^{*}$ clusters, for all points of the corresponding flow linestrings which are closer than $\varepsilon_{S}$ distance to the centroid, we snap them to the centroid. In Steps 17–18, we collate and sort all the $n_{\mathcal{F}^{*}}$ snapped linestrings $\mathcal{F}^{*}=\{(f_{1}^{*},{\bm{f}}_{1}^{*}),\dots,(f_{n_{\mathcal{F}^{*}}}^{*},{\bm{f}}_{n_{\mathcal{F}^{*}}}^{*})\}$.
@@
19:Collate snapped linestrings $\{\mathcal{F}_{1}^{*},\dots\mathcal{F}_{C^{\prime*}}^{*}\}$ into $\mathcal{F}^{*}:=\{(f_{1}^{*},{\bm{f}}_{1}^{*}),\dots,(f_{n_{\mathcal{F}^{*}}}^{*},{\bm{f}}_{n_{\mathcal{F}^{*}}}^{*})\}$
@@
A.3 Split nodes at interior intersections of linestrings
The inputs to ST_SPLITNODE are the flow linestrings $\mathcal{F}$ and the node splitting type $S$. The output are the split linestrings with added nodes $\mathcal{F}^{*}$. We deploy, for $S=$ “subdivision” in Steps 1–3, to_spatial_subdivision in the sfnetworks package (van der Meer et al., 2021), and for $S=$ “unary” in Steps 4–5, geos_unary_union in the geos package (Dunnington and Pebesma, 2020).
@@ A.3 Split nodes at interior intersections of linestrings
[…]
A.4 Blend candidate linestrings onto reference linestring
The inputs to ST_LINEBLEND are the reference linestring ${\bm{f}}_{\rm ref}$, the $n_{\mathcal{F}_{\rm cand}}$ candidate linestrings $\mathcal{F}_{\rm cand}$, and the blend tolerance $\varepsilon$. The output are the modified reference linestring and $n_{\mathcal{F}_{\rm cand}}$ modified candidate flow linestrings, all with added interior points for accurate calculation of exactly equal linestring segments. In Step 1, we initialize a local network graph $\mathcal{N}^{*}$ with the reference linestring. In Steps 2–3, we extract all points of the candidate linestrings, and use the network_blend function in the sfnetworks package (which we denote as ST_NETWORK_BLEND) to blend efficiently these points into the reference linestring (van der Meer et al., 2021). ST_NETWORK_BLEND projects the candidate linestrings onto the reference linestring, and explicitly adds them to the network, thereby creating new edges in $\mathcal{N}^{*}$. The result is a local network graph with more, shorter edges and with nodes at the projected candidate points, and whose union is the reference linestring. In Steps 4–5, we extract the $n_{\mathcal{F}_{\rm cand}}$ blended candidate linestrings with these added interior points by applying the shortest path search between the start and end point of each blended candidate linestring along the network graph $\mathcal{N}^{*}$. Step 6 involves collating the blended candidate linestrings. Step 7 is the equivalent for the blended reference linestring.
@@ A.4 Blend candidate linestrings onto reference linestring
[…]
A.5 Snap candidate-touching linestrings onto reference linestring
The inputs to ST_SNAP_CAND_TOUCH are the reference linestring ${\bm{f}}_{\rm ref}$, the $n_{\mathcal{F}_{\rm cand}}$ candidate linestrings $\mathcal{F}_{\rm cand}$, the $n_{\mathcal{F}_{\rm ct}}$ candidate-touching linestrings $\mathcal{F}_{\rm ct}$, and the snap tolerance $\varepsilon_{S}$. The output are $\mathcal{F}^{*}_{\rm ct}$ the snapped candidate-touching flow linestrings. In Steps 1–8, we iterate over each candidate-touching linestring. In Steps 2–3, we compute the intersections between the candidate-touching linestring and the boundary points of the candidate linestrings, and the respective distances. In Steps 4–7, if this intersection point is within $\varepsilon_{S}$ distance to the boundary points of ${\bm{f}}_{\rm ref}$, then we snap the candidate-touching linestring to the closest boundary point. In Step 8, if the intersection point is not within $\varepsilon_{S}$ distance, then we snap it to the nearest point on ${\bm{f}}_{\rm ref}$. These snapping operations are similar to those in Steps 15–16 in ST_SNAPNODE in Algorithm 2, and ensure that we maintain connectivity between $\mathcal{F}_{\rm ct}$ and ${\bm{f}}_{\rm ref}$. In Step 9, we collate these snapped linestrings.
Algorithm 5 ST_SNAP_CAND_TOUCH – Snap candidate-touching linestrings onto reference linestring
@@ A.5 Snap candidate-touching linestrings onto reference linestring
A.6 Compute line blending priority
@@
The inputs to ST_LINEBLEND_PRIORITY are the flow linestrings $\mathcal{F}$, the number of connected edges for the reference linestrings $k$, and the blend tolerance $\varepsilon$. The output is a set of non-overlapping reference linestrings $\mathcal{F}_{\rm ref}$, a set of unique candidate linestrings $\mathcal{F}_{\rm cand}$, and a set of (potentially repeated) candidate-touching $\mathcal{F}_{\rm ct}$ linestrings. In Steps 1–2, we set up a network graph from the linestrings $\mathcal{F}$, and then extract $\mathcal{K}$, all simple paths (i.e., all paths composed of connected, unique edges) with $k$-edges. In Steps 3–4, we compute the weighted mean of the $k$ flow values, weighted by the length of individual edges, and assign this to the entire $k$-edge path. The function ${\rm Edge}(\cdot)$ extracts the component edges from a $k$-edge path. In Step 5, we sort the flow linestrings, in descending order of their flow values and length. In Steps 6–8, we construct a maximal set of non-overlapping $k$-edge paths $\mathcal{K}^{*}$. We initialise $\mathcal{K}^{*}$ to contain the first path. We iterate through $\mathcal{K}$ and, if this path in $\mathcal{K}$ does not overlap the current $\mathcal{K}^{*}$, then we add it to $\mathcal{K}^{*}$. For $k=1$, we bypass Steps 2–8. In Steps 9–18, we step through the ${\bm{k}}_{i}^{*}\in\mathcal{K}^{*}$, starting from the linestring with the highest flow and length. In Steps 11–13, if the current flow linestring ${\bm{k}}_{i}^{*}$ has no incident edges in the candidate linestrings $\mathcal{F}_{\rm cand}$, then we set it to be a reference linestring. In Step 13, we search for candidate linestrings for ${\bm{k}}_{i}^{*}$, from those linestrings which are not already reference or candidate linestrings, according to Equation (1). In Steps 14–17, if there is at least one candidate linestring, then we update $\mathcal{F}_{{\rm ref}},\mathcal{F}_{\rm cand}$, and $\mathcal{F}_{\rm ct}$. In Step 18, we extract the candidate $\mathcal{F}_{\rm cand}$ and candidate-touching $\mathcal{F}_{\rm ct}$ linestring sets.
@@
1:Input: $\mathcal{F}$ flow linestrings, $k$ #edges, $\varepsilon$ blend tolerance
@@
A.7 Locally align segment flows using line blending
The inputs to ST_OVERLINE_LINEBLEND are the flow linestrings $\mathcal{F}$, the split node type $S$, the number of connected edges $k$, the blend tolerance $\varepsilon$, and the snap tolerance $\varepsilon_{S}$. The output are the modified reference linestring ${\bm{f}}^{*}_{\rm ref}$ and projected candidate flow linestrings $\mathcal{F}^{*}_{\rm cand}$, all with added interior points. In Step 1, we node split the flow linestrings $\mathcal{F}$. In Step 2, we node snap the flow linestrings. In Step 3, we set up the priority for the line blending, and in Step 4 we store all the linestrings that will not be modified by the line blending in $\mathcal{F}^{c}$. In Steps 5–8, we iterate over each reference linestring. In Step 6, we update the reference and candidate flow linestrings by blending the candidate linestrings. In Step 7, we apply ST_OVERLINE_PLANR within each of the $k$-edges of the updated reference linestrings. In Step 8, we compute the weighted mean flow, weighted by the length of the corresponding sub-segments, and assign it as the flow value to all sub-segments. The result from Steps 6–8 is a modified $k$-edge reference linestring with $k$ aggregated flows. In Step 9, we snap the candidate-touching linestrings to the reference linestring to maintain connectivity. In Steps 10–11, we collate the modified linestrings from Steps 5–9, with the unmodified linestrings $\mathcal{F}^{c}$ from Step 4, and sort them in descending flow and length.